### v1

# A Simultaneous ECG-PCG Acquisition System with Real-Time Burst-Adaptive Noise Cancellation

Abstract
Cardiac auscultation is an essential clinical skill, requiring excellent hearing to distinguish subtle differences in timing and pitch of heart sounds. However, diagnosing solely from these sounds is often challenging due to interference from surrounding noise, and the information may be limited. Existing solutions that adaptively cancel external noise are either not real-time or are computationally intensive, making them unsuitable for implementation in a portable system. This work proposes an end-to-end system with a real-time adaptive noise cancellation pipeline integrated into a device that simultaneously acquires electrocardiogram (ECG) and phonocardiogram (PCG) signals. The performance of the system is validated using real-world hospital noise datasets and recordings captured with the dual-modality device. For PCG and ECG signals recorded from the device in noisy hospital settings, the proposed algorithms achieved signal-to-noise ratio improvements of 37.01 dB and 30.32 dB, respectively. These results demonstrate the system’s effectiveness in enabling reliable and accessible cardiac screening, including noisy hospital environments typical of resource-constrained settings.
I Introduction
Cardiovascular diseases (CVDs) primarily arise from structural abnormalities of the heart, which may be congenital (present from birth) or acquired later in life. Congenital heart diseases affect approximately 9.41 per 1000 live births [1], causing more than 260,000 global deaths in 2017 [2], and were estimated to globally account for over 200,000 deaths in children under five in 2021 [3]. Meanwhile, CVDs in general caused an estimated 19.8 million deaths worldwide in 2022 (around 32% globally), and three-quarters of CVD deaths take place in developing (low- and middle-income) countries [4].
The stethoscope is traditionally considered the primary tool ubiquitously used for the initial assessment of cardiac conditions. However, considerable expertise is required for an accurate diagnosis using a stethoscope [5], and noise conditions surrounding clinical environments make this task further challenging [6]. Complementary to cardiac auscultation, the electrocardiogram (ECG) is an effective, non-invasive diagnostic tool that reflects the electrical activity of the heart.
Since the cyclic production of ECG and heart sounds (phonocardiogram, PCG) is biologically synchronized, the combination of these two modalities makes the localization of murmurs within the cardiac cycle as well as the identification of the first (S1) and the second (S2) heart sounds, significantly more accurate [7]. Diagnosing this using a traditional stethoscope requires considerable experience, as distinct heart murmurs, which are frequent indicators of cardiac abnormalities, can be easily missed in clinical diagnosis by a semi-trained healthcare professional [5]. Furthermore, several studies have demonstrated that combined ECG–PCG data can be used to estimate clinically meaningful parameters that reflect the heart’s contractile function and pumping capacity [8, 9, 10, 11]. Recent research on machine learning techniques also shows that incorporating and/or coupling ECG and PCG modalities can significantly improve classification accuracy between normal and pathological cardiac conditions [12, 13, 14].
Due to the advancements in sensor technologies and signal processing techniques, digital stethoscopes have gained an edge over traditional stethoscope. Today, widely used commercial digital stethoscopes can facilitate volume amplification, ambient noise reduction, and wireless transmission of sound signals between remote processing devices [15, 16, 17, 18]. However, many of these digital stethoscopes lack the functionality of synchronized ECG and PCG acquisition with wireless data transmission and real-time noise cancellation in hospital settings. Existing solutions that overcome these problems are unavailable in most developing countries, Eg: Eko Core 500[17]. In parallel, recent research has also explored wearable systems for integrated ECG–PCG monitoring [19, 5, 20] and synchronous ECG–PCG auscultation platforms [21, 22]
Despite these advancements, recording high-quality PCG signals remains challenging due to their susceptibility to ambient noise, especially in crowded or low-resource hospital environments. Such noise can distort the characteristic features of PCG signals and degrade diagnostic performance. Consequently, denoising the PCG signal is crucial before any further analysis. Prior studies have addressed PCG denoising through multistage adaptive least mean squares (LMS) filter architectures [23], deep learning–based frameworks [24], and optimization-driven adaptive filtering approaches [25]. However, an open need remains for algorithms capable of adaptively denoising PCG signals in real-time while maintaining low computational complexity for embedded implementation.
In this work, we propose a portable device for simultaneous ECG and PCG acquisition, featuring two integrated real-time pipelines. One performs ECG denoising, while the other performs adaptive noise cancellation (ANC) using a burst-adaptive normalized LMS (NLMS) filter, designed to handle sudden, non-stationary noise bursts. Finally, we quantitatively evaluate the ANC algorithm using a publicly available PCG dataset corrupted with hospital ambient noise and separately the whole system on real-world ECG-PCG recordings. This comprehensive, end-to-end system enhances the quality of both signal modalities, enabling more reliable and accessible cardiovascular diagnostics in real-world hospital settings.
Figure 1: Block diagram of the overall system architecture. Signals from dry Ag/AgCl electrodes (acquiring ECG) and digital microelectromechanical system (MEMS) microphones (acquiring PCG) are captured, processed by a microcontroller unit (MCU), and transmitted wirelessly to a computer for real-time ANC and signal visualization.
II System Overview
The functional blocks of the overall system architecture for our dual-modality device used for comprehensive cardiac screening are depicted in Fig. 1. It integrates the acquisition, processing, transmission, and denoising of two biosignals, ECG and PCG, from the body in real-time.
II-A Simultaneous Acquisition of ECG and PCG Signals
The device achieves its primary functionality of the main processing circuit (Fig. 1), using the following components: a medical-grade ADS1294 IC, with a 24-bit $\Delta\Sigma$ ADC, is used for the single-lead ECG acquisition, while CMM-4030DT digital MEMS microphones, with a sensitivity of 26-26 dB, are used for the heart and ambient sound acquisition.
Figure 2: Photographs of the prototype with the design features annotated.
A 3M Littmann Cardiology III stethoscope head is connected to a dual-microphone array to acquire PCG signals, with a 3D-printed rubber seal placed at the intersection to provide additional noise immunity. In this design, the stethoscope diaphragm provides acoustic amplification for the heart sounds before transduction, as these signals are too weak for a digital microphone to capture directly. A quad-microphone array is placed as shown in Fig. 2, with the sound ports facing outward, to effectively capture an ambient noise reference while maintaining compact device geometry.
The Lead-I ECG signal is captured using three Ag/AgCl dry electrodes, with one serving as the reference electrode to reduce common-mode noise. The placement of these electrodes is shown in Fig. 2. The MCU processes both ECG and PCG signals simultaneously, sampling ECG data at $500$ Hz and the audio data at $8$ kHz, which is subsequently downsampled to $2$ kHz. This real-time data is then streamed to the computer via a Bluetooth® Low Energy (BLE) module. The double-buffering technique is used to prevent data loss during acquisition, processing, and transmission.
[…]
Normal LMS Filter $0.1576\pm 0.3527$ $15.4595\pm 4.8141$ $0.9456\pm 0.0798$ $0.2123\pm 0.2424$
NLMS Filter $0.0259\pm 0.0237$ $19.5678\pm 4.7743$ $0.9877\pm 0.0112$ $0.1304\pm 0.0487$
Burst-Adaptive NLMS Filter $0.0229\pm 0.0201$ $20.0462\pm 4.9773$ $0.9891\pm 0.0099$ $0.1285\pm 0.0483$
Abnormal LMS Filter $0.0918\pm 0.2190$ $17.5187\pm 4.5253$ $0.9647\pm 0.0569$ $0.1541\pm 0.1782$
NLMS Filter $0.0285\pm 0.0238$ $19.2425\pm 5.2589$ $0.9863\pm 0.0119$ $0.1400\pm 0.0552$
Burst-Adaptive NLMS Filter $0.0264\pm 0.0221$ $19.5504\pm 5.3933$ $0.9873\pm 0.0112$ $0.1386\pm 0.0546$
[…]
Normal
(Bursts only) NLMS Filter $0.0546\pm 0.0487$ $17.3029\pm 3.2470$ $0.9747\pm 0.0207$ $0.1155\pm 0.0296$
Burst-Adaptive NLMS Filter $0.0327\pm 0.0231$ $19.1962\pm 3.4049$ $0.9843\pm 0.0107$ $0.1043\pm 0.0279$
Abnormal
(Bursts only) NLMS Filter $0.0496\pm 0.0400$ $17.6148\pm 3.7049$ $0.9766\pm 0.0179$ $0.1160\pm 0.0349$
Burst-Adaptive NLMS Filter $0.0316\pm 0.0194$ $19.2743\pm 3.7186$ $0.9848\pm 0.0093$ $0.1064\pm 0.0328$
II-B Real-Time Adaptive Noise Cancellation for PCG Signals
A general LMS filter uses an adaptive finite impulse response (FIR) filter for noise cancellation [26]. In this implementation, the filter is used to denoise a recorded PCG signal, $x(n)$, using a reference signal, $r(n)$, which is uncorrelated with heart sounds but closely correlated with the noise. To update the filter weights $\mathbf{w}(n)$, the LMS algorithm uses the method of steepest descent, given by
$\mathbf{w}(n+1)=\mathbf{w}(n)-\tfrac{1}{2}\mu\nabla e^{2}(n),$ (1)
where the step size $\mu$ controls the rate of convergence [26]. As an improvement, the NLMS algorithm is used when the knowledge of input signal statistics is limited [27], which is often the case in real-time filtering. It uses a step size that is normalized by the reference input power:
$\mu=\frac{\mu_{0}}{\varepsilon+\|\mathbf{r}(n)\|^{2}}.$ (2)
Here, $\mu_{0}$ is the base step size, $\|\cdot\|^{2}$ is the euclidean norm of the input vector, and $\varepsilon$ is a small positive constant.
However, in the conventional NLMS algorithm, sudden increases in the input noise power can cause the normalization term $\|\mathbf{r}(n)\|^{2}$ to become large, resulting in slow adaptation. To address this, we propose a burst-adaptive NLMS (BA-NLMS) algorithm, which dynamically scales the step size during high-energy events. In the real-time implementation, we estimate the average energy using a moving average formulation:
$\bar{E}(n)=\alpha\bar{E}(n-1)+(1-\alpha)\|\mathbf{r}(n)\|^{2},$ (3)
where $0<\alpha<1$ controls the averaging window. The effective step size $\mu_{\text{eff}}(n)$ is then adapted dynamically based on the ratio between instantaneous and average energy:
$\mu_{\text{eff}}(n)=\begin{cases}\mu_{0}\times\beta,&\text{if }\|\mathbf{r}(n)\|^{2}>\eta\bar{E}(n),\\ \mu_{0},&\text{otherwise},\end{cases}$ (4)
where $\eta$ denotes the burst detection threshold and $\beta>1$ is the step-size scaling factor. Accordingly, by simplifying the gradient vector $\nabla e^{2}(n)$ into a product of $e(n)$ and $\mathbf{r}(n)$ [26], the weight update formula in Equation 1 can be expressed as:
$\mathbf{w}(n+1)=\mathbf{w}(n)+\left(\frac{\mu_{\text{eff}}(n)}{\varepsilon+\mathbf{r}^{T}(n)\cdot\mathbf{r}(n)}\right)e(n)\mathbf{r}(n).$ (5)
II-C Real-time Denoising Pipeline for ECG Signals
The raw ECG signal is passed through three elliptic infinite impulse response filters: a $0.5$ Hz high-pass filter to eliminate baseline drift, a $150$ Hz low-pass filter to suppress high-frequency noise, and a $49.5$ - $50.5$ Hz bandstop filter to attenuate power-line interference. Elliptic filters were chosen for their sharp roll-off and high stopband attenuation at a low filter order, making them suitable for real-time processing.
[…]
Figure 4: Denoising performance of the proposed real-time filter algorithms on recorded noisy ECG and PCG signals. Spectrograms (left) and time-domain plots (right) show the improvement in the noisy ECG (input SNR of 9.08-9.08 dB) and noisy PCG (input SNR of 2.08-2.08 dB) signals, respectively.
III Results and Discussion
III-A Performance Evaluation with Artificially Induced Noise
[…]
The filter is evaluated with a length of $10$, and parameters: $\mu_{0}=0.05$, $\alpha=0.99$, $\eta=5$, $\beta=6$, and $\varepsilon=0.001$. For comparison, the standard LMS and NLMS filters were also evaluated with the same parameters. As shown in Table I, the BA-NLMS algorithm clearly outperforms both standard algorithms for all normal and abnormal PCG recordings.
To specifically evaluate performance against burst-type noise, 9 recordings containing sudden bursts were selected from our dataset and assessed. As summarized in Table II and illustrated in Fig. 3, the BA-NLMS filter demonstrated significant improvement over the standard NLMS filter. This is critical for handling non-stationary hospital noises, which are characterized by sudden amplitude changes.
III-B Prototype Performance Validation with Recorded Data
To test real-world performance, we recorded data in a noisy hospital environment and then adaptively denoised it. The last stage of our PCG denoising pipeline includes a $20$ - $200$ Hz band-pass filter, targeting the primary heart sound range [19]. Fig. 4 demonstrates the spectrograms and the time domain plots of the denoising outputs of both PCG and ECG signals, respectively. We evaluated the performance of our prototype by recording the signals simultaneously for a fixed time period of 14 seconds and calculating the SNR, using:
$\textit{SNR}=10\,\cdot\log_{10}\left[\frac{\textit{power}~{(LF)}}{\textit{power}~{(LFN)}+\textit{power}~{(HF)}}\right],$ (6)
where a low frequency range ($LF$) is expected to contain useful data, whereas a low frequency noise range ($LFN$) and a high frequency range ($HF$) are expected to contain noise. The BA-NLMS filter pipeline improved the noisy PCG signal by $\Delta\text{SNR}$ of $37.01$ dB while the noisy ECG signal was also improved, achieving a $\Delta\text{SNR}$ of $30.32$ dB.
IV Conclusion
We have proposed a portable device capable of simultaneous ECG and PCG acquisition with integrated real-time ANC. The system incorporates a novel real-time BA-NLMS algorithm that effectively handles non-stationary, burst-type noise, achieving an SNR improvement of 37.01 dB, while the real-time ECG filter achieved an SNR improvement of 30.32 dB on real-world hospital data, demonstrating the robustness of the proposed system. Future work will focus on (a) further improving the SNR performance of the prototype device, (b) implementing the denoising algorithms on the microcontroller for fully embedded operation, and (c) developing an ECG–PCG dataset recorded using the proposed device to support future diagnostic studies. Ultimately, this system has the potential to serve as a valuable tool for medical professionals to perform more reliable cardiac assessments in both clinical and low-resource environments, and expand affordable community screening for pre-existing heart diseases.
V Acknowledgements
The authors thank the staff of Lady Ridgeway Hospital for Children for providing access to record data for prototype performance evaluation and for Voex Technologies Pvt. Ltd. for funding the prototype development.